A beta^+ particle is moving at 19 500 m s^-1 paral 39.2
Answer #1

39.2]

a) F =qvB = mv2/r

so, r = mv2/qvB = [9.11*10-31 * 550*103]/[1.6*10-19 * 0.25*10-3] = 1.2cm

b) the electron will be moving counter-clockwise

c) the radius of the circular path will be the same

d) the direction of motion of the positron will be opposite to the electron, so it will be clock-wise

e) charge on alpha particle = 2e

so, v = qBr/m = [2*1.6*10-19 * 0.25*10-3 * 0.012]/[6.64*10-24] = 0.145m/s